Benzene at 20°C has a viscosity of 0.000651 Pa-s. What shear. stress is required to deform this fluid at a strain rate of 4900 s-1?




1
Expert's answer
2022-02-21T17:19:03-0500

"\mu" = 0.000651 Pa.s


"\frac{du}{dy}" = 4900s-1


= "\mu"("\frac{du}{dy}" )


= (0.000651)(4900)


=3.19Pa."x"